let the continuous random variable x denote the current measured in a thin copper wire in milliamperes. The probability density function is given by


1
Expert's answer
2021-07-13T05:10:10-0400
"f(x)=\dfrac{1}{20}, for\ 0\leq x\leq 20"

1.


"P(5<X<10)=\displaystyle\int_{5}^{10}\dfrac{1}{20}dx=[\dfrac{1}{20}x]\begin{matrix}
 10 \\
 5
\end{matrix}"

"=\dfrac{10}{20}-\dfrac{5}{20}=\dfrac{1}{4}"

2. Using Uniform distribution


"F(x)=\dfrac{x-0}{20-0}"

"P(5<X<10)=F(10)-F(5)"

"=\dfrac{10-0}{20-0}-\dfrac{5-0}{20-0}=\dfrac{1}{4}"

3.


"\mu=\dfrac{A+B}{2}=\dfrac{0+20}{2}=10"

"\sigma^2=\dfrac{(B-A)^2}{12}=\dfrac{(20-0)^2}{12}=\dfrac{100}{3}"

"\sigma=\sqrt{\sigma^2}=\dfrac{10\sqrt{3}}{3}"